# A transformational approach to collective behavior

### v2

@@ I Introduction
There have been a set of mathematical analysis methods developed, starting in the 1990’s with wavelet analysis (Mallat, 1999), that have been laying the path to the solution that we developed which is “a mathematically logical process” of renormalization, with a physical interpretation. Wavelet analysis is based on convolution of a signal with a bank of scaled wavelets. The relation of the wavelet analysis to the “singularity spectrums” of the system was recognized by Herrmann (1997) and several other authors. Wavelet decompositions were used as parameters of probability distributions that were estimated as Gaussian distributions (Strauss et al., 2003), and were estimated by Probabilistic Neural Networks (PNNs) as sums of Gaussian distributions (Glinsky et al., 2001). Wavelet analysis was extended to an iterative deep convolution by the Mallat Scattering Transformation (MST) (Mallat, 2012). Between each convolution is the nonlinear activation function, modulus. It was noted by Mallat and coworkers that the analysis became near linear if the logarithm was taken before and after the transformation (Bruna and Mallat, 2013). There have been further developments by Mallat and coworkers to include phase, in a rather Byzantine way, with Wavelet Phase Harmonics (WPH) (Mallat et al., 2020; Zhang and Mallat, 2021; Allys et al., 2020; Régaldo-Saint Blancard et al., 2023); and to relate it to Wilson renormalization with Wavelet Conditional Renormalization Group (WC-RG) (Marchand et al., 2022).
@@
Note there has been a large Lagrangian detour taken from this canonical route down the road of regrouping the Mayer Cluster Expansion, based on the weakness of correlation (that is, the BBGKY hierarchy (Nicholson, 1983)), and the weakness of coupling (that is, quantum perturbation theory by Feynman (Feynman, 1961) and others). This has lead to the super convergence of the Mayer Cluster Expansion being reduced to only asymptotic convergence, and to the infinities that appear in the Wilson renormalization procedure. The connection of the MST to the S-matrix (to second order) and to renormalization was first recognized by Glinsky (2014).
@@
Addressing the two issues with Quantum Field Theory highlighted by Dirac, inspired by Dirac’s view that group symmetry is primal, and building upon the work of Gell-Mann, Heisenberg, Mallat, and genAI, we develop this canonical transformation approach to collective behavior. In Sec. II, we start by assuming a group symmetry to the motion of the collective, leading to the conservation of a real function, $H(p,q)$, that is the infinitesimal generator of a Lie group, $\mathscr{H}$. The finite generator of the canonical transformation, $S_{P}(q)$, is derived from the infinitesimal generator, $H(p,q)$, by the solution of the Hamilton-Jacobi equation. The motion of the collective field, $f(x)$, is then generated by a canonical generating functional, $S_{p}[f(x)]$, a deep deconvolution with the complex logarithm (that is, $\ln(z)=\ln|z|+\text{i}\,\arg(z)$) as the activation function. The formula for this generating functional, the Heisenberg Scattering Transformation (HST), is derived in Sec. III, and shown in Eq. (21). This gives the Taylor expansion of $S_{P}(q)$, or the functional Taylor expansion of $S_{p}[f(x)]$, or the S-matrix, or the Mayer Cluster expansion, or the Wigner-Weyl Transformation of the collective field, $f(x)$. In other words, $S^{m}_{i}=d^{m}H(\beta)/d\beta^{m}_{i}=\left|\beta_{i}(z)\right>$, where $\left|\beta_{i}(z)\right>$ are the principal components of the HST. The theory is quantized in Sec. IV. How to fit, simulate, and control (that is, ponderomotively stabilize) the collective system is described in Sec. V. Section VI shows how to couple one collective system to another collective system forming systems-of-systems. The practical computer implementation of the HST is outlined in Sec. VII. Finally, there is a discussion of the results and conclusions in Sec. VIII.
@@ II Canonical Generating Functional Approach
The four elemental forces are associated with the four elemental symmetries, shown in Fig. 1. In order of increasing scale, they are: SU(3) for the strong force, SU(2) for the weak force, SU(1) for the EM force, and $\mathbb{H}_{g}$ for the gravitational force. The same is true, given any Hamiltonian function, $H(p,q)$. For a pendulum clock with the Hamiltonian $H_{p}(p,q)=p^{2}/2-\omega^{2}\,\cos(q)$, one can form the complex Lie group $\mathbb{H}_{p}=\mathscr{H}_{p}\otimes\text{Ad}(\mathscr{H}_{p})$. For the weather with the Hamiltonian $H_{w}(p,q)$, one can form the complex Lie group $\mathbb{H}_{w}=\mathscr{H}_{w}\otimes\text{Ad}(\mathscr{H}_{w})$. Both of these collective systems are intermediate in scale between the EM force and the gravitational force.
@@
The finite generating functional form of the action, $S_{p}[f(x)]$, is the Heisenberg Scattering Transformation (HST). The form of this transformation is derived in Sec. III and appears as Eq. (21). A graphical representation of the HST expansion is shown in Fig. 4. These expansion coefficients can be viewed several ways. They can be viewed as the Mayer Cluster Expansion in the order of the correlation. For the Mayer Cluster Expansion, $S_{1}$ tells how one body is distributed, $S_{2}$ tells how two bodies are correlated, $S_{3}$ tells how three bodies are correlated, $S_{4}$ tells how four bodies are correlated, and $S_{m}$ tells how $m$ bodies are correlated. They can be viewed as the S-matrix or the $m$-body Scattering Cross Section expansion in the order of the scattering. For Heisenberg’s S-matrix, $S_{1}$ is the one body scattering cross section, $S_{2}$ is the two body scattering cross section, $S_{3}$ is the three body scattering cross section, $S_{4}$ is the four body scattering cross section, and $S_{m}$ is the $m$-body scattering cross section. They also can be viewed as the $m$-body Green’s Function expansion, and the complex $m^{\text{th}}$-order curvatures of the manifold.
@@
The functional Taylor expansion coefficients of the HST, $S_{p}[f(x)]$, are equal to the Taylor expansion coefficients of $S_{P}(q)$. The results of a Principal Components Analysis (PCA) on the output of the HST, $\left|\beta_{i}(z)\right>$, where there is one component for each field, are the functional Taylor expansion coefficients of the HST or the S-matrix. They are
[…]
@@ III Theory of the Heisenberg Scattering Transformation (HST)
The transformation we endeavour to construct takes the dynamics to a manifold that is a linear subspace of a Hilbert space with basis vectors that are the solution to the Renormalization Group Equations (RGEs). In this subspace, the dynamics is geodesic motion with the topology given by the Hamiltonian function $H(z)$, an analytic function on $\mathbb{C}^{n}$, where $n$ is the number of fields. Another way of looking at this is that the motion is harmonic or holomorphic. We start by constructing a logarithmic generator of the function $\bar{H}\equiv\text{i}H$. First, Taylor expand the function $\bar{H}$ about $z_{0}$ giving
$\bar{H}(z)=\sum_{m=0}^{\infty}{\frac{1}{m!}S_{m}(z_{0})\,(z-z_{0})^{m}},$ (6)
@@
$S_{m}(z)=\frac{d^{m}\bar{H}(z)}{dz^{m}}=\frac{dH_{m}}{dz},$ (7)
which serves as a definition of $H_{m}(z)$. Now, we need to find an expression for $H_{m}$ given $H_{m-1}$. Start by assuming that $H_{m}$ is a functional of the field and the field momentum, that is
$H_{m}(z(x))=F[\pi(x),f(x)].$ (8)
@@
$\begin{split}dH_{m}(z)&=\text{i}\frac{dz_{k}}{|z|}=\text{i}\left(\frac{dz_{\parallel}}{|z|}+\text{i}\frac{dz_{\perp}}{|z|}\right)\\ &=\text{i}\left(\frac{dJ}{J}+\text{i}\theta\right)=\text{i}\,d(\ln(J\text{e}^{\text{i}\theta}))\\ &=\text{i}\,d(\ln z)=d(\text{i}\,\ln z).\end{split}$ (13)
@@
$R_{0}(z)\equiv z+\text{e}^{\text{i}\arg(z)},$ (14)
one gets the recursion relation
$\begin{split}H_{m}&=\text{i}\,\ln z=\text{i}\,\ln(R_{0}(z_{k}))\\ &=\text{i}\,\ln R_{0}\,\psi_{k_{m}}\star H_{m-1}.\end{split}$ (15)
@@
$\ln(R_{0}(z))\xrightarrow[|z|\to 0]{}z$ (17)
is a compact mapping. We use $f(x)$ as shorthand for any $F[\pi(x),f(x)]$. Start with
$H_{0}=\text{i}\,\ln R_{0}\,f(x).$ (18)
@@
$\boxed{H_{m}[f(x)](z)=\phi_{kx}\star\left(\prod_{n=1}^{m}{\text{i}\ln R_{0}\psi_{k_{n}}\star}\right)\text{i}\ln R_{0}f(x),}$ (21)
where $z\equiv k+\text{i}\,x$. This Logarithmic Generating functional (LG) generates, by construction, $\text{i}H(z)=\bar{H}(z)$. It is not the exponential generator, propagator, or partition function, given by $U=\text{e}^{-\text{i}H\tau}$ or $Z=\text{e}^{\text{i}S}$, as is done in the Lagrangian approach. It follows that
@@
which means that the transformation has flattened the space so that the infinitesimal generator, $\text{i}H$, is now the finite group generator. We have transformed to the basis $\left|\beta(z)\right>$ where $\beta\equiv p+\text{i}\,q$, the Hilbert space of all possible solutions to the RGEs, corresponding to $S_{m}=dH_{m}/dz$, the S-matrix. The motion will be confined to an $n$-dimensional complex linear hyperplane $\mathbb{C}^{n}$, where $n$ is the number of fields, and will be geodesic motion on that submanifold, determined by the analytic function $H(\beta)$, with $m^{\text{th}}$-order complex curvature
$\frac{d^{m}H(\beta)}{d\beta^{m}}=\frac{dH_{m}}{dz}=S_{m}\sim\text{i}\,d(\ln H_{m})\sim\frac{1}{\beta_{\text{rc},m}},$ (23)
where $\beta_{\text{rc},m}$ is the $m^{\text{th}}$-order complex radius of curvature. Streamlines of the motion generated by $H$ (the $\mathscr{H}$ group action) are given by $\text{Re}(H(\beta))=\text{constant}$, and streamlines of $dH$ (the $\text{Ad}(\mathscr{H})$ group action) are given by $\text{Im}(H(\beta))=\text{constant}$.
@@
Now lets review what the HST is doing. The $\psi_{k}\star$ is generating $z_{k}(x)$ a parametric trajectory of the analytic function $\text{i}H$. The $\ln R_{0}$ conformal (canonical) transformation is flattening the space onto the cylinder by transforming into polar coordinates about $R_{0}$. The i conformal transformation is rotating $\pi/2$ from the tangent direction to the gradient (covector) direction. Another perspective views it as a canonical transformation that is exchanging the field momentum with the field. The choice of $R_{0}$ ensures that the repeated application of the mapping converges to a circle of radius $\pi$ about the origin on the cylinder. This is because $R_{0}$ is a fixed point of $\ln$, that is $\ln(R_{0}(z))\to z$ as $|z|\to 0$. The convergence to the origin is exponential for large deviations, then logarithmic for small deviations.
@@
$\begin{split}\ln(R_{0}(z))&\xrightarrow[|z|\to 0]{}z\\ &\xrightarrow[|z|\to\infty]{}\ln|z|\quad\text{(Glinsky and Maupin)}\\ &\xrightarrow[|z|\to\infty]{}\ln|z|\,\text{e}^{\text{i}\,\arg(z)/\ln|z|}\\ &\sim\sum_{k=0}^{n}{|z|\,\text{e}^{\text{i}\,k\,\arg{(z)}}}\quad\text{(that is, WPH)}\\ &\xrightarrow[n=0]{}|z|\quad\text{(that is, MST)}.\end{split}$ (24)
@@
The process of one iteration of the HST can be viewed in this simple way. The convolution generates an analytic trajectory on the complex plane. The ln changes to polar coordinates after translating the origin to the fixed point of the mapping, all canonical and analytic transformations. This flattens the space, and exposes the curvature of the manifold as the expansion coefficients, $d^{m}H/d\beta^{m}$. This process is simply taking the canonical derivative, $\text{i}\,d(\ln)$, where the exterior derivative is taken by $\psi_{k}\star$.
@@
The relationship to the Werner Heisenberg’s 1943 S-matrix ideas comes from the fact that the S-matrix or curvature is $d^{m}H/d\beta^{m}$, so we are calculating the S-matrix by calculating $H(\beta)$. Just take the derivatives. This is why Eq. (21) is called the Heisenberg Scattering Transformation. Second, the HST is also closely related to the ideas of Eugene Wigner and Hermann Weyl, when they constructed the Wigner-Weyl transformation. They wanted to transform from the collective field domain to $\mathbb{R}^{n}$, then solve an ODE. The HST follows this basic philosophy by transforming from the collective field domain to $\mathbb{C}^{n}$, then solving Laplace’s Equation for the minimal or geodesic surface.
The concept of the S-matrix was first introduced by Werner Heisenberg in 1943 (Heisenberg, 1943). He lost interest in it, most likely, because he lacked concepts of analyticity. It was picked up again by Chew and Low (1956) in 1955 and was most completely developed by Lev Landau (Landau, 1959) and Richard Cutkosky (Cutkosky, 1960) by 1960. A good survey of this work was written by Chew (1961). This work was based on the concepts of analyticity and unitarity. It further used the expansion of an exponential generator employed by Richard Feynman in the path integral approach to quantum field theory. Furthermore, it evaluated the resulting integrals using a Fourier basis. It had no physical basis for the analyticity. Because of this, the theory was incomplete and did not form a well defined theoretical structure. It was also very difficult to calculate, and was not evaluated past second order. This line of research was abandoned in favor of the path integral formulation.
@@
This work is also closely related to the Wigner-Weyl transformation (Case, 2008; Wigner, 1932; Weyl, 1950). They wanted to transform from the collective field domain to $\mathbb{R}^{n}$, then solve an ODE. The problems were they: (1) used a global Fourier basis not compatible with calculation on a manifold, (2) only calculated to second order, and (3) did not include the logarithmic transform. This resulted in complicated corrections to the commutator, divergences in the evaluation, and an incomplete transformation.

### v6

@@ I Introduction
There have been a set of mathematical analysis methods developed, starting in the 1990’s with wavelet analysis (Mallat, 1999), that have been laying the path to the solution that we developed which is “a mathematically logical process” of renormalization, with a physical interpretation. Wavelet analysis is based on convolution of a signal with a bank of scaled wavelets. The relation of the wavelet analysis to the “singularity spectrums” of the system was recognized by Herrmann (1997) and several other authors. Wavelet decompositions and their moments were used as parameters of probability distributions that were estimated as Gaussian distributions (Strauss et al., 2003; Glinsky et al., 2003), and were estimated by Probabilistic Neural Networks (PNNs) as sums of Gaussian distributions (Glinsky et al., 2001). Wavelet analysis and its moments was extended and formalized to an iterative deep convolution by the Mallat Scattering Transformation (MST) (Mallat, 2012). Between each convolution is the nonlinear activation function, modulus. It was noted by Mallat and coworkers that the analysis became near linear if the logarithm was taken before and after the transformation (Bruna and Mallat, 2013). There have been further developments by Mallat and coworkers to include phase, in a rather Byzantine way, with Wavelet Phase Harmonics (WPH) (Mallat et al., 2020; Zhang and Mallat, 2021; Allys et al., 2020; Régaldo-Saint Blancard et al., 2023); and to relate it to Wilson renormalization with Wavelet Conditional Renormalization Group (WC-RG) (Marchand et al., 2022).
@@
Note there has been a large Lagrangian detour taken from this canonical route down the road of regrouping the Mayer Cluster Expansion, based on the weakness of correlation (that is, the BBGKY hierarchy (Nicholson, 1983; Glinsky, 2024a)), and the weakness of coupling (that is, quantum perturbation theory by Feynman (Feynman, 1961) and others). This has lead to the super convergence of the Mayer Cluster Expansion being reduced to only asymptotic convergence, and to the infinities that appear in the Wilson renormalization procedure. The connection of the MST to the S-matrix (to second order) and to renormalization was first recognized by Glinsky (2014).
@@
Addressing the two issues with Quantum Field Theory highlighted by Dirac, inspired by Dirac’s view that group symmetry is primal, and building upon the work of Gell-Mann, Heisenberg, Mallat, and genAI, we develop this canonical transformation approach to collective behavior. In Sec. II, we start by assuming a group symmetry to the motion of the collective, leading to the conservation of a real function, $H(p,q)$, that is the infinitesimal generator of a Lie group, $\mathscr{H}$. The finite generator of the canonical transformation, $S_{P}(q)$, is derived from the infinitesimal generator, $H(p,q)$, by the solution of the Hamilton-Jacobi equation. The motion of the collective field, $f(x)$, is then generated by a canonical generating functional, $S_{p}[f(x)]$, a deep deconvolution with the complex logarithm (that is, $\ln(z)=\ln|z|+\text{i}\,\arg(z)$) as the activation function. The formula for this generating functional, the Heisenberg Scattering Transformation (HST), is derived in Sec. III, and shown in Eq. (21). This gives the Taylor expansion of $S_{P}(q)$, or the functional Taylor expansion of $S_{p}[f(x)]$, or the S-matrix, or the Mayer Cluster expansion, or the Wigner-Weyl Transformation of the collective field, $f(x)$. The theory is quantized in Sec. IV. How to fit, simulate, and control (that is, ponderomotively stabilize) the collective system is described in Sec. V. Section VI shows how to couple one collective system to another collective system forming systems-of-systems. The practical computer implementation of the HST is outlined in Sec. VII. Finally, there is a discussion of the results and conclusions in Sec. VIII.
@@ II Canonical Generating Functional Approach
The four elemental forces are associated with the four elemental symmetries, shown in Fig. 1. In order of increasing scale, they are: SU(3) for the strong force, SU(2) for the weak force, SU(1) for the EM force, and $\mathbb{H}_{g}$ for the gravitational force. The same is true, given any Hamiltonian function, $H(p,q)$. For a pendulum clock with the Hamiltonian $H_{p}(p,q)=p^{2}/2-\omega^{2}\,\cos(q)$, one can form the complex Lie group $\mathbb{H}_{p}=\mathscr{H}_{p}\otimes\text{Ad}(\mathscr{H}_{p})$. For the weather with the Hamiltonian $H_{w}(p,q)$, one can form the complex Lie group $\mathbb{H}_{w}=\mathscr{H}_{w}\otimes\text{Ad}(\mathscr{H}_{w})$. A well known three dimensional weather model is the Lorenz model (Lorenz, 1963). Both of these collective systems are intermediate in scale between the EM force and the gravitational force.
@@
The finite generating functional form of the action, $S_{p}[f(x)]$, is the Heisenberg Scattering Transformation (HST). The form of this transformation is derived in Sec. III and appears as Eq. (21). A graphical representation of the HST expansion is shown in Fig. 4. These expansion coefficients can be viewed several ways. They can be viewed as the Mayer Cluster Expansion in the order of the correlation. For the Mayer Cluster Expansion, $S_{1}$ tells how one body is distributed, $S_{2}$ tells how two bodies are correlated, $S_{3}$ tells how three bodies are correlated, $S_{4}$ tells how four bodies are correlated, and $S_{m}$ tells how $m$ bodies are correlated. They can be viewed as the S-matrix or the $m$-body Scattering Cross Section expansion in the order of the scattering. For Heisenberg’s S-matrix, $S_{1}$ is the one body scattering cross section, $S_{2}$ is the two body scattering cross section, $S_{3}$ is the three body scattering cross section, $S_{4}$ is the four body scattering cross section, and $S_{m}$ is the $m$-body scattering cross section. They also can be viewed as the $m$-body Green’s Function expansion, and the $m^{\text{th}}$-order complex curvatures of the manifold.
@@
The functional Taylor expansion coefficients of the HST, $S_{p}[f(x)]$, are equal to the Taylor expansion coefficients of $S_{P}(q)$. The results of a Principal Components Analysis (PCA) on the output of the HST, $\left|\beta_{i}(z)\right>$, where there is one component for each field, are the functional Taylor expansion coefficients of the HST or the S-matrix, $S_{m}$. They are
$\begin{split}\frac{\delta^{m}S[f(x)]}{\delta f^{m}}&=\frac{d^{m}S(z(x))}{dz^{m}}=S_{m}\\ &=\left|\beta(z)\right>=\frac{d^{m}S(\beta)}{d\beta^{m}}=\frac{d^{m}S(q)}{dq^{m}},\end{split}$ (3)
@@ III Theory of the Heisenberg Scattering Transformation (HST)
The transformation we endeavor to construct takes the dynamics to a manifold that is a linear subspace of a Hilbert space with basis vectors that are the solution to the Renormalization Group Equations (RGEs). In this subspace, the dynamics is geodesic motion with the topology given by the Hamiltonian function $H(z)$, an analytic function on $\mathbb{C}^{n}$, where $n$ is the number of fields. Another way of looking at this is that the motion is harmonic or holomorphic. We start by constructing a logarithmic generator of the function $\bar{H}\equiv\text{i}H$. First, Taylor expand the function $\bar{H}$ about $z_{0}$ giving
$\bar{H}(z)=\sum_{m=0}^{\infty}{\frac{1}{m!}H_{m}(z_{0})\,(z-z_{0})^{m}},$ (6)
@@
$H_{m}(z)=\frac{d^{m}\bar{H}(z)}{dz^{m}}=\frac{dS_{m}(z)}{dz},$ (7)
which serves as a definition of $S_{m}(z)$. Now, we need to find an expression for $S_{m}$ given $S_{m-1}$. Start by assuming that $S_{m}$ is a functional of the field and the field momentum or the complex field, that is
$S_{m}(z(x))=F[\pi(x),f(x)]=F[f(x)=\pi(x)+\text{i}\,f(x)].$ (8)
@@
$\begin{split}dS_{m}(z)&=\text{i}\frac{dz_{k}}{|z|}=\text{i}\left(\frac{dz_{\parallel}}{|z|}+\text{i}\frac{dz_{\perp}}{|z|}\right)\\ &=\text{i}\left(\frac{dJ}{J}+\text{i}\theta\right)=\text{i}\,d(\ln(J\text{e}^{\text{i}\theta}))\\ &=\text{i}\,d(\ln z)=d(\text{i}\,\ln z).\end{split}$ (13)
@@
$R_{0}(z)\equiv\frac{1}{\text{i}}h^{-1}(2z/\pi),$ (14)
and $h(z)\equiv(z+1/z)/2$, one gets the recursion relation
$\begin{split}S_{m}&=\text{i}\,\ln z=\text{i}\,\ln(R_{0}(z_{k}))\\ &=\text{i}\,\ln R_{0}\,\psi_{k_{m}}\star S_{m-1}.\end{split}$ (15)
@@
$\text{i}\,\ln(R_{0}(z))\xrightarrow[\epsilon\to 0,\,\text{for }z\in(-\pi/2+\text{i}\epsilon,\pi/2+\text{i}\epsilon)]{}z$ (17)
is a compact mapping for $z\in(-\pi/2,\pi/2)$. We use $f(x)$ as shorthand for any $F[\pi(x),f(x)]$. Start with
$S_{0}=\text{i}\,\ln R_{0}\,f(x).$ (18)
@@
$\boxed{S_{m}[f(x)](z)=\phi_{kx}\star\left(\prod_{n=1}^{m}{\text{i}\ln R_{0}\psi_{k_{n}}\star}\right)\text{i}\ln R_{0}f(x),}$ (21)
[…]
@@
[…]
$\frac{d^{m-1}H(\beta)}{d\beta^{m-1}}=\frac{dS_{m-1}}{dz}=S_{m}\sim\frac{d(\ln z)}{dz}\sim\frac{1}{\beta_{\text{rc},m}},$ (23)
where $\beta_{\text{rc},m}$ is the $m^{\text{th}}$-order complex radius of curvature, and we used the expression $S_{m-1}\sim\ln z$ given by Eq. (13). Streamlines of the motion generated by $H$ (the $\mathscr{H}$ group action) are given by $\text{Re}(H(\beta))=\text{constant}$, and streamlines of $dH$ (the $\text{Ad}(\mathscr{H})$ group action) are given by $\text{Im}(H(\beta))=\text{constant}$.
@@
We take a closer look at the analytic rectifier function, $R(z)\equiv\text{i}\,\ln(R_{0}(z))$. Both the $R(z)$ and $R_{0}(z)$ conformal mappings are shown in Fig. 6. Note that the $R_{0}(z)$ mapping takes the contour around the branch cut between $-\pi/2$ and $\pi/2$ to a contour around the unit circle. There are two sheets to the Riemann surface of $R_{0}$. The first sheet is mapped to the area outside the unit circle, the other sheet is mapped to the area inside the unit circle. The complete conformal $R(z)$ mapping takes the contour twice circulating around the branch cut between $-\pi/2$ and $\pi/2$, through the branch cut for $\exp(z)$, to the contour that circulates once around $-\pi$ and $\pi$. The part of the contour between $-\pi/2+\text{i}\epsilon$ and $\pi/2+\text{i}\epsilon$ is mapped to itself. Remember that the Riemann surface for $\exp(z)$ has an infinite number of sheets. The mapping is compact, converging to the real axis. The result is a compact conformal mapping converging to the fixed contour of a unit helix unrolled around the real axis, with repeated application and bination.
Figure 6: The conformal mappings, $R_{0}(z)$ and $R(z)$. Branch points are shown as the black dots, and branch cuts are shown as the thin gray rectangles. Reference points on the contours are shown as colored dots. Black arrows indicate convergence of the compact mapping, $\ln(z)$. Note that the part of the contour between $-\pi/2+\text{i}\epsilon$ and $\pi/2+\text{i}\epsilon$ is mapped to itself, by the analytic rectifier function, $R(z)$. The result is a compact conformal mapping converging to the fixed contour of a unit helix unrolled around the real axis, with repeated application of $R(z)$ and bination.
Now lets review what the HST is doing. The $\psi_{k}\star$ is generating $z_{k}(x)$ a parametric trajectory of the analytic function $\text{i}H$. The $\ln R_{0}$ conformal (canonical) transformation is flattening the space onto the cylinder by transforming into polar coordinates about $R_{0}$. The i conformal transformation is rotating $\pi/2$ from the tangent direction to the gradient (covector) direction. Another perspective views it as a canonical transformation that is exchanging the field momentum with the field. This can be recognized by the fact that $\pi(x)+\text{i}\,f(x)$ and $f(x)+\text{i}\,\pi(x)$ are canonically conjugate fields. The choice of $R_{0}$ ensures that the repeated application of the mapping converges to a circle of radius $\pi$ about the origin on the cylinder. This is because $R_{0}$ is a fixed point of $\ln$, that is $\ln(R_{0}(z))\to z$ as $|z|\to 0$. The convergence to the origin is exponential for large deviations, then logarithmic for small deviations.
@@
$\begin{split}\ln(R_{0}(z))&\xrightarrow[z\to(-\pi/2,\pi/2)]{}z/\text{i}\\ &\xrightarrow[|z|\to\infty]{}\ln|z|\quad\text{(Glinsky and Maupin)}\\ &\xrightarrow[|z|\to\infty]{}\ln|z|\,\text{e}^{\text{i}\,\arg(z)/\ln|z|}\\ &\sim\sum_{k=0}^{n}{|z|\,\text{e}^{\text{i}\,k\,\arg{(z)}}}\quad\text{(that is, WPH)}\\ &\xrightarrow[n=0]{}|z|\quad\text{(that is, MST)}.\end{split}$ (24)
@@
The process of one iteration of the HST can be viewed in this simple way. The convolution generates an analytic trajectory on the complex plane. The ln changes to polar coordinates after translating the origin to the fixed point of the mapping, all canonical and analytic transformations. This flattens the space, and exposes the curvature of the manifold as the expansion coefficients or S-matrix, $S_{m}$. This process is simply taking the canonical derivative, $\text{i}\,d(\ln)$, where the exterior derivative is taken by $\psi_{k}\star$.
@@
The relationship to the Werner Heisenberg’s 1943 S-matrix ideas comes from the fact that the S-matrix is the curvature, $S_{m}=d^{m-1}H/d\beta^{m-1}$, so we are calculating the expansion coefficients of $H(\beta)$ by calculating the S-matrix. Just take the derivative, $H_{m}=dS_{m}/dz$. This is why Eq. (21) is called the Heisenberg Scattering Transformation. Second, the HST is also closely related to the ideas of Eugene Wigner and Hermann Weyl, when they constructed the Wigner-Weyl transformation. They wanted to transform from the collective field domain to $\mathbb{R}^{n}$, then solve an ODE. The HST follows this basic philosophy by transforming from the collective field domain to $\mathbb{C}^{n}$, then solving Laplace’s Equation for the minimal or geodesic surface.
The concept of the S-matrix was first introduced by Werner Heisenberg in 1943 (Heisenberg, 1943). He lost interest in it, most likely, because he lacked concepts of analyticity. It was picked up again by Chew and Low (1956) in 1955 and was most completely developed by Lev Landau (Landau, 1959) and Richard Cutkosky (Cutkosky, 1960) by 1960. A good survey of this work was written by Chew (1961). This work was based on the concepts of analyticity and unitarity. It further used the perturbative expansion (in terms of the strength of the field coupling constant) of an exponential generator employed by Richard Feynman in the path integral approach to quantum field theory. Furthermore, it evaluated the resulting integrals using a Fourier basis. It had no physical basis for the analyticity. Because of this, the theory was incomplete and did not form a well defined theoretical structure. It was also very difficult to calculate, and was not evaluated past second order. This line of research was abandoned in favor of the path integral formulation.
@@
This work is also closely related to the Wigner-Weyl transformation (Case, 2008; Wigner, 1932; Weyl, 1950). They wanted to transform from the collective field domain to $\mathbb{R}^{n}$, then solve an ODE. The problems were they: (1) used a global Fourier basis not compatible with calculation on a manifold, (2) only calculated to second order, and (3) did not include the logarithmic transform. This resulted in complicated corrections to the commutator, divergences in the evaluation, and an incomplete transformation. Another way of looking at this is that Wigner and Weyl constructed a transformation from the Hilbert space to $\mathbb{R}^{n}$, but found that they could not guarantee that the sympletic structure would be preserved. This is because sympletic structure is not built into the algebra of $\mathbb{R}^{n}$, but symplectic structure is built into $\mathbb{C}^{n}$. The Cauchy-Riemann equations are Hamilton’s equations.